What is the Mass of 6.10^23 atoms of oxygen
1
Expert's answer
2020-06-30T07:47:17-0400

We need to find amount of substance "n=\frac{N}{N_A}" and then multiply it by oxygen molar mass

"m = \frac{N}{N_A}M = \frac{6\times10^{23}}{6.02\times10^{23}mol^{-1}}\times32\frac{g}{mol} \approx 31.89g"

(What Avogadro constant value do you use? If it's 6*10^23 the answer would be 32g)

Answer: 31.89g